# Citadel: Context Similarity Based Deep Learning Framework Bug Finding

### v4

@@ Abstract.
With the application of deep learning technology, tools of DL framework testing are in high demand. Existing DL framework testing tools have limited coverage of bug types. For example, they lack the capability of finding performance bugs, which are critical for DL models regarding performance, economics, and the environment. Moreover, existing tools are inefficient, generating hundreds of test cases with few trigger bugs. In this paper, we propose Citadel, a method that accelerates bug finding in terms of efficiency and effectiveness. We observe that many DL framework bugs are similar due to the similarity of operators and algorithms belonging to the same family. Orthogonal to existing bug-finding tools, Citadel aims to find new bugs that are similar to reported ones that have known test oracles. Citadel defines context similarity to measure the similarity of DL framework API pairs and automatically generates test cases with oracles for APIs that are similar to the problematic APIs in existing bug reports. Citadel effectively detects 77 and 74 API bugs on PyTorch and TensorFlow, many of which, e.g., 13 performance bugs, cannot be detected by existing tools. Moreover, 35.40% of test cases generated by Citadel can trigger bugs significantly transcending the state-of-the-art method (3.90%).
Deep Learning Testing, Library Testing, Software Testing, Deep Learning
ccs: Software and its engineering Software libraries and repositories
@@ 1. Introduction
With the development of Deep Learning (DL) techniques, DL-powered systems are playing an increasingly significant role in software development. For example, Microsoft has developed a new search engine powered by DL techniques to enhance the search results (mic, 2023). Moreover, the global AI software market is expected to increase from \$138 billion in 2022 to \$1,094 billion by 2032 (mar, 2022). As the backbone of DL-powered systems, DL frameworks (e.g., TensorFlow and PyTorch) empower developers by offering API functions to create, train, optimize, and deploy DL-powered systems. These frameworks support diverse domains, providing societal benefits in areas like image recognition (He et al., 2016), self-driving (Grigorescu et al., 2020), and natural language processing (Li, 2018). Similar to traditional software systems, DL frameworks can also have bugs, which can lead to erroneous outputs, increased system overhead, and even crashes for DL-powered systems, thereby jeopardizing user property and personal safety (Blanco, 2023), and contributing to energy inefficiency and environmental issues (Nistor et al., 2015; Jin et al., 2012; Lacoste et al., 2019). Consequently, there is a pressing need for tools capable of identifying bugs in DL frameworks.
@@
Despite these advancements, existing DL testing tools have notable limitations. Firstly, existing testing tools have limited coverage of bug types. For example, they fail to detect performance bugs that can significantly impact DL model training and inference speed and degrade responsiveness, resulting in energy waste and environmental concerns (Nistor et al., 2015; Jin et al., 2012; Lacoste et al., 2019; Bryant et al., 2003), especially for large DL models like GPT-3 (Patterson et al., 2021). The performance bug shown in Fig. 2 causes the time overhead to increase to 2.33 times its original value and a substantial carbon footprint. However, current testing methods cannot detect such performance bugs in DL API functions. Secondly, existing bug-finding tools exhibit inefficiencies in generating test cases that trigger bugs. These tools often leverage random walks or heuristic algorithms to generate test cases. However, due to the huge search space of API arguments and inputs, such approaches often generate numerous test cases but only a small fraction of them trigger actual bugs. For example, DeepREL generates an excess of 330,000 test cases, yet only 1.23% of them have the potential to trigger bugs. Requiring hundreds of test cases to detect a single bug makes current DL framework testing tools very inefficient in detecting bugs.
@@
Building upon this observation, we propose Citadel, a tool that accelerates the finding of bugs in terms of efficiency and effectiveness. Orthogonal to existing tools, Citadel aims to uncover new bugs that are similar to reported ones that have known test oracle, regardless of bug types. It leverages reported bugs on one API function to create test cases for its analogous API functions, effectively addressing the aforementioned limitations observed in existing work, and can easily co-work with other testing methods to accelerate bug discovery. Compared with existing tools which can only detect status bugs and value bugs, Citadel has better bug type coverage and effectiveness. It has the capability to detect bugs regardless of their types, such as performance bugs caused by errors in the underlying implementation or optimization, including unexpected time or memory overhead. Moreover, Citadel is more effective and efficient in test case generation. It adopts the code that has triggered a bug on a problematic API function to create test cases for its analogous API functions. Essentially, it leverages prior knowledge rather than heuristics to explore potential API bugs in the new context, significantly improving the chances of finding bugs. To be specific, we first collect existing bug reports and identify problematic APIs. Citadel then utilizes both static and dynamic analyses on DL framework source code and unit test cases for the identification of analogous API functions. In this process, it extracts context information (e.g., APIs’ call stacks) to gauge the similarity between API functions, a concept referred to as context similarity in Citadel. For a collected problematic API, Citadel modifies the bug-triggering code from its bug report to generate new test cases for its analogous API functions. Throughout this process, Citadel addresses two potential differences between the API functions: differences in arguments and dimensions, if they exist. Finally, Citadel executes the generated test cases, employing the buggy behavior of the problematic API function as a test oracle to effectively identify potential new bugs in the target API function.
Our evaluation demonstrates that Citadel successfully identified a total of 77 API bugs in PyTorch and 74 in TensorFlow, including 58 and 66 previously unreported bugs, of which 36 and 56 have been confirmed. Additionally, 49 of these bugs are detected by analogous API pairs that existing approaches do not cover. Furthermore, a noteworthy 35.40% of the test cases synthesized by Citadel expose bugs, significantly surpassing the 0.74%, 1.23%, and 3.90% bug-triggering capacity exhibited by the test cases generated by DocTer, DeepREL, and TitanFuzz respectively.
@@
  • We develop a novel test case generation method for DL frameworks that leverages the knowledge from confirmed API bugs to synthesize new test cases and uncover new bugs in analogous API functions, regardless of bug types.
  • We develop a prototype Citadel based on the proposed idea. The experimental results on PyTorch and TensorFlow show that Citadel respectively detects 77 and 74 API bugs, and 36 and 56 of them have been confirmed or fixed by developers after reporting. 35.40% of test cases generated by Citadel can be used to trigger bugs.
@@ 2.1. DL Framework API Functions and Models
DL Framework APIs. Like traditional software programs, DL frameworks use various API functions to call source code functions and perform operations. Taking PyTorch (Paszke et al., 2019) as an example, its API functions include performing basic matrix operations (e.g., torch.mul for multiply operation), calculating loss functions (e.g., torch.nn.MSELoss for measuring mean squared error), and building models layers (e.g., torch.nn.Conv2d for convolution layers). When calling an API function, users first need to assign values for its required and optional arguments, where the values of required arguments are mandatory to provide and the optional arguments have their default values in APIs. Then, the API function runs the underlying source code that performs corresponding calculations and operations on the hardware (e.g., CPU and GPU) and obtains tensors, Boolean values, etc. as the result.
DL Models. A DL model is a parameterized function $F_{\theta}:X\mapsto Y$, where $x\in X$ is an $m-$dimensional input and $y\in Y$ is the corresponding output label. Typically, a DL model is composed of several connected layers, and an $n$-layered model can be represented as $F_{\theta}=l_{1}\circ l_{2}\circ\cdot\cdot\cdot\circ l_{n}$, where $l$ represents a layer and $\theta$ is the model weight. Each layer $l_{i}$ in the model can be constructed by several DL framework API functions. Therefore, a DL model also be represented as a directed acyclic graph (DAG) in that API functions are nodes, and the returned values of API (e.g., tensors) are edges. Running and training a DL model $F$ on the input-output pairs ($x_{i},y_{i}$) is essentially calling a series of API functions and passing their outputs based on the topological sorting of its computation graph (Deng et al., 2022).
@@ 2.2. DL Framework Testing
DL framework testing methods construct test cases (e.g., models) to explore potential behaviors of DL frameworks and discover bugs. Depending on the generated test cases, existing DL framework testing methods can be mainly divided into model-level testing and API-level testing (Wei et al., 2022; Deng et al., 2022).
Model-level testing. These testing methods usually build a large number of models and apply mutation strategies on models to explore the potential bugs of the APIs and layers in the model. To construct test oracles, prior work performs differential testing by building and testing the same model on multiple DL frameworks (Guo et al., 2020; Pham et al., 2019; Gu et al., 2022; Wang et al., 2020b). CRADLE is one of the first to use this method to test DL framework bugs. Based on Keras (Ketkar and Ketkar, 2017) which can build and train models on different DL frameworks as backends, it conducts differential testing on three frameworks (i.e., TensorFlow, CNTK, and Theano) and finds 12 bugs. Additionally, Muffin (Gu et al., 2022) creatively designs the data tracking method to apply differential testing on the training phase of models and finally discovers 39 new bugs. Although the model-level methods obtain outstanding test results, they still have great limitations in applications. Due to limitations imposed by the test model, these methods typically support only a limited subset of API functions related to the models. For instance, existing research (Wei et al., 2022) reports that LEMON covers only 35 TensorFlow APIs. Furthermore, since the test oracle relies on the implementations of multiple frameworks, inconsistencies detected during testing are often difficult to verify as real bugs, which hampers the effectiveness of the bug detection process (Guo et al., 2020; Pham et al., 2019).
@@
DL framework bugs. DL framework bugs can be mainly divided into three types through symptoms, i.e., status, value, and performance bugs (Chen et al., 2022). Status bugs affect the execution status of DL API and model, including crashes, segmentation faults, exceptions, etc. Value bugs that are caused by numerical errors in the computation of DL operators include inconsistent outputs and NaN (Not A Number) outputs. Existing framework testing tools focus on the above two types of bugs (Guo et al., 2020; Deng et al., 2022; Xie et al., 2022). Performance bugs refer to those caused by errors in the underlying implementation or optimization, including unexpected time or memory overhead.
@@ 2.3. Code Similarity Measurement
Code similarity measurement aims to evaluate the similarity of multiple code blocks and find out the potential code clone, plagiarism, and refactoring. Existing static approaches proposed methods based on the metrics, texts, and tokens (Faidhi and Robinson, 1987; Luo et al., 2014; Misu and Sakib, 2017; Wang et al., 2023). Researchers also measure code similarity based on Abstract Syntax Trees (ASTs) and graphs (e.g., control flow graphs (CFG)) (Zhang et al., 2019; Chae et al., 2013). In addition, some research proposes the functional similarity between programs from the perspectives of input and output and function calls (Su et al., 2016; McMillan et al., 2012), etc. However, these methods and tools are usually designed for code snippets in one single programming language, but DL framework API functions execute on both Python and C++ source code and involve various wrappers, which poses a challenge in evaluating the similarity between these API functions. Inspired by existing approaches, Citadel defines and calculates the context similarity to match and test DL framework API functions. Citadel calculates the similarity of source code blocks from the perspectives of inputs, outputs, and functions. Then it combines the above results with the called functions and traces of different DL framework APIs to match APIs that have a similar functionality and execution context.
@@ 3. Motivation
$\bullet$ Existing approaches have limited coverage on bug types. They focus on status and value bugs, lacking the capability of detecting others, e.g., performance bugs. Existing methods typically utilize the output results of the same or equivalent API function on different frameworks/devices to construct pseudo test oracles to identify status bugs (e.g., crashes) and value bugs (e.g., NaN outputs) (Wang et al., 2022; Gu et al., 2022; Deng et al., 2022; Guo et al., 2020; Pham et al., 2019; Deng et al., 2023). However, they cannot construct test oracles to detect performance bugs due to the difficulty of obtaining test oracles. Consequently, the detection of performance bugs still heavily relies on manual identification by developers.
@@ 4. Design
We observe that DL framework API functions naturally fall into groups, and API functions in a group (e.g., Conv1d, Conv2d, and Conv3d in PyTorch) execute similar operators and algorithms and have similar functionality and implementations, despite potential differences in their arguments (e.g., different dimensions). Moreover, through sampling and analysis of over 300 real bug reports and corresponding patches from DL frameworks, we find that API functions within the same groups are susceptible to similar bugs due to the same erroneous implementation of underlying functions (tor, 2022d, b, 2021). As such, Citadel is designed to exploit the similarity between DL framework API functions to identify bugs.
@@
Fig. 1 shows the overview of Citadel. The inputs to Citadel include DL framework repositories, real-world bug cases collected from the framework issues, and the signature similarity between DL API functions (Deng et al., 2022). Specifically, in this paper, Citadel is applied to the PyTorch and TensorFlow repositories, two of the most widely used DL frameworks, which have garnered 87K and 189K stars on GitHub, respectively. Bug cases are collected from the issue lists of these two frameworks, including reproducible buggy code and problematic API function (§ 4.1). The workflow of Citadel begins by extracting context information from the DL framework’s repository, including the API function call stacks and analogous function groups derived from the framework’s source code. In this process, the dynamic profiler (§ 4.3) generates unit test cases for DL API functions and records their call stacks, capturing the source code functions invoked during execution. Simultaneously, the static analyzer (§ 4.4) examines the DL framework’s source code and clusters analogous functions based on argument and callee similarity. Citadel proposes context similarity, which utilizes API call stacks to measure the similarity between API functions. Additionally, the analogous source code functions are treated as identical during similarity calculation. Leveraging the context similarity and the signature similarity (Deng et al., 2022), the API matcher (§ 4.5) matches analogous API pairs. Furthermore, the matcher verifies the arguments of each API pair and discards those with unsolvable argument mismatches. For a problematic API function (i.e., source API) and its analogous API function (i.e., target API), the test case generator utilizes the reproducible buggy code of the source API, which is collected in the bug case list, to synthesize new test cases for the target API (§ 4.6). The test case evaluator then executes new test cases and leverages the buggy behavior exhibited by the source API to identify new bugs in the target API, including status, value, and performance bugs (§ 4.7). Finally, Citadel reports the newly detected API bugs to the user.
@@
Here we provide a real-world performance bug found by Citadel as an example (Lines 1-4 in  Fig. 2(a)) to show how it works. One collected bug on Conv2d (Fig. 2(b)) is reported saying that the ‘groups’ argument in this function fails to speed up the training and inference. Grouped convolution aims to employ multiple kernels and produce multiple channel outputs to increase the network efficiency (Krizhevsky et al., 2017; Gong et al., 2021). Therefore, the group convolution is anticipated to bring a lower time overhead compared to executing these convolution layers independently. The code in the collected bug uses the execution time of independent convolution layers to estimate an upper bound on expected time overhead of group convolution, which is accepted by the developers, and finds that the actual overhead of group convolution (11.79 s) is much greater than the expected upper bound (8.91 s), therefore identified the performance bug on Conv2d. Citadel analyzes the context information (e.g., call stacks) of DL API functions to construct pairs of analogous API functions that share context similarity. One such pair consists of Conv2d and LazyConvTranspose2d. Then, Citadel generates test cases for LazyConvTranspose2d based on the reproducible code of the problematic API Conv2d. For each analogous API pair, Citadel analyzes the arguments of the two API functions to identify differences and makes adjustments to the buggy code accordingly to construct a test case for the target, ensuring the generated test case is executable. In this instance, Citadel remove ‘in_channels’ to resolve the difference between APIs’ arguments, which is highlighted by green ( 1). In addition, Citadel leverages the method in the source bug report of Conv2d to estimate the expected upper bound on the time overhead of the grouped LazyConvTranspose2d layers using the overhead of independent LazyConvTranspose2d layers. The new test case reveals that LazyConvTranspose2d with ‘group’ argument also exhibits a higher time overhead than executing these layers individually, which is the same anomalous behavior as the reported bug in Conv2d ( 2). Specifically, as shown in Fig. 2(a), when we set the ‘out_channels’ to 512 and ‘groups’ to 4 in LazyConvTranspose2d layer and construct model_1 with eight such group convolution layers, the actual time cost of training model_1 is 46.75 s, which is markedly higher than the time cost of executing these layers individually (20.09 s). Finally, Citadel discovers a performance bug on the LazyConvTranspose2d API function, which has been confirmed by developers (per, 2023a).
Fig. 2. Performance Bug on LazyConvTranspose2d

### v5

@@ Abstract.
With the application of deep learning technology, tools of DL framework testing are in high demand. Existing DL framework testing tools have limited coverage of bug types. For example, they lack the capability of effectively finding performance bugs, which are critical for DL models regarding performance, economics, and the environment. Moreover, existing tools are inefficient, generating hundreds of test cases with few trigger bugs. In this paper, we propose Citadel, a method that accelerates bug finding in terms of efficiency and effectiveness. We observe that many DL framework bugs are similar due to the similarity of operators and algorithms belonging to the same family. Orthogonal to existing bug-finding tools, Citadel aims to find new bugs that are similar to reported ones that have known test oracles. Citadel defines context similarity to measure the similarity of DL framework API pairs and automatically generates test cases with oracles for APIs that are similar to the problematic APIs in existing bug reports. Citadel effectively detects 58 and 66 API bugs on PyTorch and TensorFlow (excluding those rejected by developers or duplicates of prior reports), many of which, e.g., 13 performance bugs, cannot be detected by existing tools. Moreover, 35.40% of test cases generated by Citadel can trigger bugs significantly transcending the state-of-the-art method (3.90%).
Deep Learning Testing, Deep Learning Library, Software Testing
copyright: acmlicensedjournal: TOSEMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3773993ccs: Software and its engineering Software libraries and repositories
@@ 1. Introduction
With the development of Deep Learning (DL) techniques, DL-powered systems are playing an increasingly significant role in software development. For example, Microsoft has developed a new search engine powered by DL techniques to enhance the search results (mic, 2023). Moreover, the global AI software market is forecasted to increase from \$257 billion in 2025 to \$1,459 billion by 2034 (mar, 2025). As the backbone of DL-powered systems, DL frameworks (e.g., TensorFlow and PyTorch) empower developers by offering API functions to create, train, optimize, and deploy DL-powered systems. These frameworks support diverse domains, providing societal benefits in areas like image recognition (He et al., 2016), self-driving (Grigorescu et al., 2020), and natural language processing (Li, 2018). Similar to traditional software systems, DL frameworks can also have bugs, which can lead to erroneous outputs, increased system overhead, and even crashes for DL-powered systems, thereby jeopardizing user property and personal safety (Blanco, 2023), and contributing to energy inefficiency and environmental issues (Nistor et al., 2015; Jin et al., 2012; Lacoste et al., 2019). Consequently, there is a pressing need for tools capable of identifying bugs in DL frameworks.
@@
Despite these advancements, existing DL testing tools have notable limitations. Firstly, existing testing tools have limited coverage of bug types. For example, they can hardly detect performance bugs that can significantly impact DL model training and inference speed and degrade responsiveness, resulting in energy waste and environmental concerns (Nistor et al., 2015; Jin et al., 2012; Lacoste et al., 2019; Bryant et al., 2003), especially for large DL models like GPT-3 (Patterson et al., 2021). The performance bug shown in Fig. 2 causes the time overhead to increase to 2.33 times its original value and a substantial carbon footprint. However, current testing methods cannot detect such performance bugs in DL API functions. Secondly, existing bug-finding tools exhibit inefficiencies in generating test cases that trigger bugs. These tools often leverage random walks or heuristic algorithms to generate test cases. However, due to the huge search space of API arguments and inputs, such approaches often generate numerous test cases, but only a small fraction of them trigger actual bugs. For example, DeepREL generates an excess of 330,000 test cases, yet only 1.23% of them have the potential to trigger bugs. Requiring hundreds of test cases to detect a single bug makes the current DL framework testing tools very inefficient in detecting bugs.
@@
Building upon this observation, we propose Citadel, a tool that accelerates the finding of bugs in terms of efficiency and effectiveness. Orthogonal to existing tools that explore new anomalous behaviors and report bugs, Citadel aims to uncover new bugs that are similar to reported ones that have known test oracles, regardless of bug types. It leverages reported bugs on one API function to create test cases for its analogous API functions, effectively addressing the aforementioned limitations observed in existing work, and can easily co-work with other testing methods to accelerate bug discovery. Compared with existing tools, which can only detect status bugs and value bugs, Citadel has better bug type coverage and effectiveness. It has the capability to detect bugs regardless of their types, such as performance bugs caused by errors in the underlying implementation or optimization, including unexpected time or memory overhead. Moreover, Citadel is more effective and efficient in test case generation. It adopts the code that has triggered a bug on a problematic API function to create test cases for its analogous API functions. Essentially, it leverages prior knowledge rather than heuristics to explore potential API bugs in the new context, significantly improving the chances of finding bugs. To be specific, we first collect existing bug reports and identify problematic APIs. Citadel then utilizes both static and dynamic analyses on the DL framework source code and unit test cases for the identification of analogous API functions. In this process, it extracts context information (e.g., APIs’ call stacks) to gauge the similarity between API functions, a concept referred to as context similarity in Citadel. For a collected problematic API, Citadel modifies the bug-triggering code from its bug report to generate new test cases for its analogous API functions. Throughout this process, Citadel addresses two potential differences between the API functions: differences in arguments and dimensions, if they exist. Finally, Citadel executes the generated test cases, employing the buggy behavior of the problematic API function as a test oracle to effectively identify potential bugs in the target API function.
Our evaluation demonstrates that Citadel successfully identified a total of 77 API bugs in PyTorch and 74 in TensorFlow, including 58 and 66 previously unreported bugs, of which 36 and 56 have been confirmed. Additionally, 49 of these bugs are detected by analogous API pairs that existing approaches do not cover. Furthermore, a noteworthy 35.40% of the test cases synthesized by Citadel expose bugs, significantly surpassing the 0.74%, 1.23%, and 3.90% bug-triggering capacity exhibited by the test cases generated by DocTer, DeepREL, and TitanFuzz, respectively.
@@
  • We develop a novel test case generation method for DL frameworks that leverages the knowledge from confirmed API bugs to synthesize new test cases and uncover bugs in analogous API functions, regardless of bug types.
  • We develop a prototype Citadel based on the proposed idea. The experimental results on PyTorch and TensorFlow show that Citadel detects 58 and 66 previously unknown API bugs, respectively, among which 36 and 56 have been confirmed or fixed by developers after reporting. 35.40% of test cases generated by Citadel can be used to trigger bugs.
@@ 2.1. DL Framework API Functions and Models
DL Framework APIs. Like traditional software programs, DL frameworks use various API functions to call source code functions and perform operations. Taking PyTorch (Paszke et al., 2019) as an example, its API functions include performing basic matrix operations (e.g., torch.mul for multiply operation), calculating loss functions (e.g., torch.nn.MSELoss for measuring mean squared error), and building model layers (e.g., torch.nn.Conv2d for convolution layers). When calling an API function, users first need to assign values for its required and optional arguments, where the values of required arguments are mandatory to provide, and the optional arguments have their default values in APIs. Then, the API function runs the underlying source code that performs corresponding calculations and operations on the hardware (e.g., CPU and GPU) and obtains tensors, Boolean values, etc. as the result.
DL Models. A DL model is a parameterized function $F_{\theta}:X\mapsto Y$, where $x\in X$ is an $m$-dimensional input and $y\in Y$ is the corresponding output label. Typically, a DL model is composed of several connected layers, and an $n$-layered model can be represented as $F_{\theta}=l_{1}\circ l_{2}\circ\cdot\cdot\cdot\circ l_{n}$, where $l$ represents a layer and $\theta$ is the model weight. Each layer $l_{i}$ in the model can be constructed by several DL framework API functions. Therefore, a DL model can also be represented as a directed acyclic graph (DAG) in that API functions are nodes, and the returned values of API (e.g., tensors) are edges. Running and training a DL model $F$ on the input-output pairs ($x_{i},y_{i}$) is essentially calling a series of API functions and passing their outputs based on the topological sorting of its computation graph (Deng et al., 2022).
@@ 2.2. DL Framework Testing
DL framework testing methods construct test cases (e.g., models) to explore abnormal behaviors of DL frameworks and discover bugs Depending on the generated test cases, existing DL framework testing methods can be mainly divided into model-level testing and API-level testing (Wei et al., 2022; Deng et al., 2022).
Model-level testing. These testing methods usually build a large number of models and apply mutation strategies on models to explore the potential bugs of the APIs and layers in the model. To construct test oracles, prior work performs differential testing by building and testing the same model on multiple DL frameworks (Guo et al., 2020; Pham et al., 2019; Gu et al., 2022; Wang et al., 2020b). CRADLE is one of the first to use this method to test DL framework bugs. Based on Keras (Ketkar and Ketkar, 2017), which can build and train models on different DL frameworks as backends, it conducts differential testing on three frameworks (i.e., TensorFlow, CNTK, and Theano) and finds 12 bugs. Additionally, Muffin (Gu et al., 2022) creatively designs the data tracking method to apply differential testing on the training phase of models and finally discovers 39 new bugs. Although the model-level methods obtain outstanding test results, they still have great limitations in applications. Due to limitations imposed by the test model, these methods typically support only a limited subset of API functions related to the models. For instance, existing research (Wei et al., 2022) reports that LEMON covers only 35 TensorFlow APIs. Furthermore, since the test oracle relies on the implementations of multiple frameworks, inconsistencies detected during testing are often difficult to verify as real bugs, which hampers the effectiveness of the bug detection process (Guo et al., 2020; Pham et al., 2019).
@@
DL framework bugs. DL framework bugs can be mainly divided into three types through symptoms, i.e., status, value, and performance bugs (Chen et al., 2022). Status bugs affect the execution status of the DL API and model, including crashes, segmentation faults, exceptions, etc. Value bugs that are caused by numerical errors in the computation of DL operators include inconsistent outputs and NaN (Not A Number) outputs. Existing framework testing tools focus on the above two types of bugs (Guo et al., 2020; Deng et al., 2022; Xie et al., 2022). Performance bugs refer to those caused by errors in the underlying implementation or optimization, including unexpected time or memory overhead.
@@ 2.3. Code Similarity Measurement
Code similarity measurement aims to evaluate the similarity of multiple code blocks and find potential code clones, plagiarism, and refactoring. Existing static approaches proposed methods based on the metrics, texts, and tokens (Faidhi and Robinson, 1987; Luo et al., 2014; Misu and Sakib, 2017; Wang et al., 2023). Researchers also measure code similarity based on Abstract Syntax Trees (ASTs) and graphs (e.g., control flow graphs (CFG)) (Zhang et al., 2019; Chae et al., 2013). In addition, some research proposes the functional similarity between programs from the perspectives of input and output and function calls (Su et al., 2016; McMillan et al., 2012), etc. However, these methods and tools are usually designed for code snippets in one single programming language, but DL framework API functions execute on both Python and C++ source code and involve various wrappers, which poses a challenge in evaluating the similarity between these API functions. Inspired by existing approaches, Citadel defines and calculates the context similarity to match and test DL framework API functions. Citadel calculates the similarity of source code blocks from the perspectives of inputs, outputs, and functions. Then it combines the above results with the called functions and traces of different DL framework APIs to match APIs that have a similar functionality and execution context.
@@ 3. Motivation
$\bullet$ Existing approaches have limited coverage on bug types. They focus on status and value bugs, lacking the capability of effectively detecting others, e.g., performance bugs. Existing methods typically utilize the differential testing techniques to construct pseudo test oracles for bug detection. These approaches compare the output results of the same or equivalent API functions on different frameworks/devices to identify status bugs (e.g., crashes) and value bugs (e.g., NaN outputs) (Wang et al., 2022; Gu et al., 2022; Deng et al., 2022; Guo et al., 2020; Pham et al., 2019; Deng et al., 2023). However, they are generally unable to construct test oracles to detect performance bugs due to the difficulty of obtaining test oracles. In addition, although some metamorphic testing methods have successfully identified several performance bugs (Wei et al., 2022), their effectiveness is constrained by the manually designed metamorphic relations. As a result, they can only test the specific API behaviors (e.g., those related to tensor types) and are unable to identify broader categories of performance bugs (e.g., the LazyConvTranspose2d bug in Fig. 2).
@@ 4. Design
We observe that DL framework API functions naturally fall into groups. For example, in PyTorch, while APIs like Conv1d, Conv2d, and Conv3d expose parameter signatures designed to accommodate different data dimensions, they call similar or even the same operators and algorithms during runtime, indicating that their functionality and implementation are highly similar. Existing research has confirmed that this phenomenon of code cloning is prevalent in DL frameworks (Assi et al., 2025). Code cloning not merely significantly increases software maintenance costs but also accelerates the propagation of defects, that is, implementation errors in one API are highly likely to exist in other analogous API functions within the same group (Islam et al., 2016). To verify this hypothesis, we sample and manually analyze over 300 real bug reports and corresponding patches from DL frameworks and find that API functions within the same group are highly susceptible to similar or even the same implementation errors, leading to a series of bugs across multiple APIs (tor, 2022d, b, 2021). For example, prior work (Xie et al., 2022) has detected a series of crash bugs on the conv1d, conv2d, and conv3d APIs of PyTorch, which actually have the same implementation error (i.e., missing checks on the variable groups) and can eventually be fixed by the same patch [^1] https://github.com/pytorch/pytorch/pull/77919. Based on the above findings, this paper proposes CITADEL, which aims to systematically exploit the similarities in functionality and implementation between APIs (i.e., contextual similarity) and combine them with known bug reports to efficiently discover bugs in DL framework APIs.
@@
Fig. 1 shows the overview of Citadel. The inputs to Citadel include DL framework repositories, real-world bug cases collected from the framework issues, and the signature similarity between DL API functions (Deng et al., 2022). Specifically, in this paper, Citadel is applied to the PyTorch and TensorFlow repositories, two of the most widely used DL frameworks, which have garnered 87K and 189K stars on GitHub, respectively. Bug cases are collected from the issue lists of these two frameworks, including reproducible buggy code and problematic API functions (§ 4.1). The workflow of Citadel begins by extracting context information from the DL framework’s repository, including the API function call stacks and analogous function groups derived from the framework’s source code. In this process, the dynamic profiler (§ 4.3) generates unit test cases for DL API functions and records their call stacks, capturing the source code functions invoked during execution. Simultaneously, the static analyzer (§ 4.4) examines the DL framework’s source code and clusters analogous functions based on argument and callee similarity. Citadel proposes context similarity, which utilizes API call stacks to measure the similarity between API functions. Additionally, the analogous source code functions are treated as identical during similarity calculation. Leveraging the context similarity and the signature similarity (Deng et al., 2022), the API matcher (§ 4.5) matches analogous API pairs. Furthermore, the matcher verifies the arguments of each API pair and discards those with unsolvable argument mismatches. For a problematic API function (i.e., source API) and its analogous API function (i.e., target API), the test case generator utilizes the reproducible buggy code of the source API, which is collected in the bug case list, to synthesize new test cases for the target API (§ 4.6). The test case evaluator then executes new test cases and leverages the buggy behavior exhibited by the source API to identify bugs in the target API, including status, value, and performance bugs (§ 4.7). Finally, Citadel reports the newly detected API bugs to the user.
Fig. 2. Performance Bug on LazyConvTranspose2d
@@
Here we provide a real-world performance bug found by Citadel as an example (Lines 1-4 in  Fig. 2(a)) to show how it works. One collected bug on Conv2d (Fig. 2(b)) reports that the ‘groups’ argument in this function fails to speed up the training and inference. Grouped convolution aims to employ multiple kernels and produce multiple channel outputs to increase the network efficiency (Krizhevsky et al., 2017; Gong et al., 2021). Therefore, the group convolution is anticipated to bring a lower time overhead compared to executing these convolution layers independently. The code in the collected bug uses the execution time of independent convolution layers to estimate an upper bound on expected time overhead of group convolution, which is accepted by the developers, and finds that the actual overhead of group convolution (11.79 s) is much greater than the expected upper bound (8.91 s), therefore identified the performance bug on Conv2d. Citadel analyzes the context information (e.g., call stacks) of DL API functions to construct pairs of analogous API functions that share context similarity. One such pair consists of Conv2d and LazyConvTranspose2d. Then, Citadel generates test cases for LazyConvTranspose2d based on the reproducible code of the problematic API Conv2d. For each analogous API pair, Citadel analyzes the arguments of the two API functions to identify differences and makes adjustments to the buggy code accordingly to construct a test case for the target, ensuring the generated test case is executable. In this instance, Citadel remove ‘in_channels’ to resolve the difference between APIs’ arguments, which is highlighted by green ( 1). In addition, Citadel leverages the method in the source bug report of Conv2d to estimate the expected upper bound on the time overhead of the grouped LazyConvTranspose2d layers using the overhead of independent LazyConvTranspose2d layers. The new test case reveals that LazyConvTranspose2d with ‘group’ argument also exhibits a higher time overhead than executing these layers individually, which is the same anomalous behavior as the reported bug in Conv2d ( 2). Specifically, as shown in Fig. 2(a), when we set the ‘out_channels’ to 512 and ‘groups’ to 4 in LazyConvTranspose2d layer and construct model_1 with eight such group convolution layers, the actual time cost of training model_1 is 46.75 s, which is markedly higher than the time cost of executing these layers individually (20.09 s). Finally, Citadel discovers a performance bug on the LazyConvTranspose2d API function, which has been confirmed by developers (per, 2023a).